OpenStudy (anonymous):

Rewrite with a rational exponent 5 radical 20

OpenStudy (anonymous):

Is your question:\[\sqrt[5]{20}\]

OpenStudy (anonymous):

or \[5\sqrt{20}\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

That's okay. \[\sqrt[4]{a}=a^{1/4}\]. So what do you think is your answer?

OpenStudy (anonymous):

20 1/5?

OpenStudy (anonymous):

20^(1/5)... yes :).

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

=-D